OpenStudy (anonymous):

@jim_thompson5910 I got 1575 for the area of the rectangle, but then I'm stuck .

jimthompson5910 (jim_thompson5910):

which rectangle? the bigger one? or the smaller one?

OpenStudy (anonymous):

The bigger one, I did L = 35 W= 45

jimthompson5910 (jim_thompson5910):

what is the area of the smaller one

OpenStudy (anonymous):

That's what I'm trying to figure out

jimthompson5910 (jim_thompson5910):

it's 35 ft by 25 ft

OpenStudy (anonymous):

I was just about to ask that! :D

jimthompson5910 (jim_thompson5910):

notice how the markings don't go all the way across the larger rectangle

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

1575-875 = __ ?

jimthompson5910 (jim_thompson5910):

you should have 45*55 = 2,475 for the larger rectangle sorry I just noticed that error

jimthompson5910 (jim_thompson5910):

so it's 2,475 - 875 = ??

OpenStudy (anonymous):

Oh I see! 1600

jimthompson5910 (jim_thompson5910):

correct